OpenStudy (anonymous):

What is the slope of the line? A) -3 B) - 1 3 C) 1 3 D) 3

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

No, do you know what the formula for slope is?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Thats correct, in this circumstance though an easier one to use is Rise over Run

OpenStudy (anonymous):

So what is your rise?

OpenStudy (anonymous):

Is it going vertical?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

And run is horizontal, so it would be a fraction of the vertical amount over the horizontal amount, so do you have an idea as to what it is now?

OpenStudy (anonymous):

uhh im going with 3

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

would that be my answer?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

thanks alot!

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

Hey can you help with one more? @Wriddler

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

4 + x 7 = 2 A) -14 B) 10 C) 12 D) 42

OpenStudy (anonymous):

4+x/7=2

TheSmartOne (thesmartone):

\[\huge\color{blue}{Welcome~To~OpenStudy!}\] @katt345272 @~ If you have another question then make a new question and post it! Happy OpenStudying, TheSmartOne

OpenStudy (anonymous):

Yeah xD

OpenStudy (anonymous):

Close this question and make a new thread